OpenStudy (anonymous):

I think my brain has vapor lock. I can't seem to solve for y. Please help X=y/(y-1)

OpenStudy (anonymous):

@precal

OpenStudy (phi):

multiply both sides by (y-1) as a first step

OpenStudy (phi):

what do you get?

OpenStudy (anonymous):

Oops it is supposed to be (y+1) I get X(y+1)=y

OpenStudy (phi):

yes now distribute the x (multiply x* y and x*1)

OpenStudy (anonymous):

Xy+1x=y

OpenStudy (phi):

btw 1x is usually just written as x

OpenStudy (phi):

then subtract y from both sides, and while you are at it, subtract x from both sides

OpenStudy (anonymous):

This is where I am confused am I moving the single y over to the left so to speak?

OpenStudy (anonymous):

And I thought you would have to dived both sides by x?

OpenStudy (phi):

xy + x= y you can do "the same thing" to both sides, and it stays equal: xy + x -y = y - y simplify y-y to be 0 xy + x -y =0 now do the same for x: subtract x from both sides

OpenStudy (phi):

and yes, the reason for this is to "move" y to the left and x to the right side

OpenStudy (anonymous):

After subtracting x from both sides Xy-y=-x

OpenStudy (phi):

now "factor out" y from the left side. This is the opposite of "distributing"

OpenStudy (anonymous):

Y(x-1)=-x

OpenStudy (phi):

now divide both sides by (x-1)

OpenStudy (anonymous):

I guess then divide both sides by x-1?

OpenStudy (anonymous):

Y=-x/(x-1) But the book answer is Y=x/(x-1)

OpenStudy (phi):

yes. You have the correct answer, and I was going to point out one other thing. \[ y= \frac{-x}{(x-1)} =\frac{-x}{(x-1)}\cdot \frac{-1}{-1} = \frac{x}{-(x-1)}= \frac{x}{1-x}\]

OpenStudy (phi):

notice that -(x-1) is the same as -1*(x-1) distribute the -1 to get -1*x + -1*-1 which is -x +1 rewrite -x+1 ast 1-x

OpenStudy (phi):

if the book says x/(x-1) it looks like a typo somewheres....

OpenStudy (anonymous):

I was originally trying to find the inverse of F(x)=x/(x+1)

OpenStudy (phi):

We can test your inverse. pick x=3 (some arbitrary number) F(x)= F(3)= 3/4 if we call the inverse of F(x) G then G(3/4) should give back 3 G(x) = -x/(x-1) G(3/4)= -3/4 / (3/4-4/4) = -3/4 / -1/4 = 3 so it works.

OpenStudy (anonymous):

The graph of my answer does not look like the graph of the book answer. And I never could find three points (whole numbers) to verify. I only found two but they work with both my answer and the book.

OpenStudy (phi):

do they graph the original function? if the original function was F(x)= x/(x-1) then its inverse would be G(x)= x/(x-1)

OpenStudy (anonymous):

Original function F(x)=x/(x+1) I graphed it using a TI

OpenStudy (precal):